Alex had Y toy cars. He gave 11 of them to his friend. Later his dad bought him twice as many toy cars as he had in the beginning. How many toy cars did Alex have in the beginning of this story if now he has 55 of them?

Answer by Okamiden(22)

He starts with Y toy cars. Y is an unknown number of cars that we'll have to find out.
.
Start : Y
.
He gives away 11.
.
Y - 11
.
His dad gives him twice as many as he had in the beginning. In other words, he gives him twice the amount of what he started with. He started with Y, so his dad gives him 2*y. We add this 2Y on top of the Y - 11 he currently has :
.
Y - 11 + 2Y
.
Finally, we're told this final amount adds up to 55 toys.
.
Y - 11 + 2Y = 55.
.
That's our equation, we need to solve it to find Y :
.
Y - 11 + 2Y = 55
.
We group the Y and 2Y together :
3Y - 11 = 55
.
We add +11 on both sides:
3Y = 55 +11
.
3Y = 66
.
Y = 22
So the original amount of toys was 22.
.
Hope that helps!